# Inversion of Magnetic Data using Learned Dictionaries and Scale Space

### v1

@@ Abstract
Magnetic data inversion is an important tool in geophysics, used to infer subsurface magnetic susceptibility distributions from surface magnetic field measurements. This inverse problem is inherently ill-posed, characterized by non-unique solutions, depth ambiguity, and sensitivity to noise. Traditional inversion approaches rely on predefined regularization techniques to stabilize solutions, limiting their adaptability to complex or diverse geological scenarios. In this study, we propose an approach that integrates variable dictionary learning and scale-space methods to address these challenges. Our method employs learned dictionaries, allowing for adaptive representation of complex subsurface features that are difficult to capture with predefined bases. Additionally, we extend classical variational inversion by incorporating multi-scale representations through a scale-space framework, enabling the progressive introduction of structural detail while mitigating overfitting. We implement both fixed and dynamic dictionary learning techniques, with the latter introducing iteration-dependent dictionaries for enhanced flexibility. Using a synthetic dataset to simulate geological scenarios, we demonstrate significant improvements in reconstruction accuracy and robustness compared to conventional variational and dictionary-based methods. Our results highlight the potential of learned dictionaries, especially when coupled with scale-space dynamics, to improve model recovery and noise handling. These findings underscore the promise of our data-driven approach for advance magnetic data inversion and its applications in geophysical exploration, environmental assessment, and mineral prospecting.
@@ 1 Introduction
However, magnetic inversion is a classic example of a highly ill-posed inverse problem, characterized by the non-uniqueness of solutions, where multiple subsurface models can equally explain the same observed magnetic anomalies, complicating the determination of a definitive solution [27]. Depth ambiguity further compounds the problem, as magnetic signals attenuate with increasing depth, thereby limiting the resolution of deeper structures [27]. Noise in the data and variability in magnetization properties, such as induced and remanent components, can significantly distort the inversion outcomes [11]. To address these challenges, classical inversion methods incorporate regularization, such as $\ell_{1}$-$\ell_{2}$ norms, to balance sparsity and smoothness [15, 27]. Probabilistic frameworks, including linear stochastic inversion introduced by Franklin [12] and refined by Tarantola and Valette [23, 19], initially provided a foundation for addressing inversion ambiguities. Building on this probabilistic approach, Bosch et al. [3, 4] used Monte Carlo methods for gravity inversion to derive posterior probability densities, capturing the range of plausible models consistent with the data. Chasseriau and Chouteau [7] emphasized integrating prior geological information through three-dimensional gravity data inversion constrained by an a priori covariance model. Incorporating spatial correlations of subsurface properties enhances stability and geological realism, overcoming the limitations of traditional deterministic methods. Nonetheless, these conventional approaches rely on pre-determined regularization schemes, which may limit their adaptability to complex geological scenarios.
In this work, we propose a data-driven approach to magnetic inversion by integrating dictionary learning [13] with a scale-space framework. Our framework aims to learn a dynamic dictionary that evolves over iterations, creating a versatile inversion process that can approximate complex models which fixed dictionaries often fail to represent. Numerical experiments demonstrate that our approach not only outperforms classical methods in terms of accuracy and robustness but also offers greater flexibility in handling diverse geological structures. This analysis highlights the promise of a scale-space approach as an alternative for standard dictionary learning in geophysical inversion.
@@ 2 Overview of the problem
Magnetic inversion aims to reconstruct a three-dimensional magnetic susceptibility model from two-dimensional surface magnetic data. This problem is inherently ill-posed, as it is under-determined, exhibits limited depth resolution, and is subject to the non-uniqueness of potential fields. According to Gauss’s theorem, infinitely many source distributions can produce the same observed magnetic field. To address these challenges, inversion methods incorporate prior information and regularization strategies to ensure physically plausible and geologically meaningful solutions.
This section delves into the mathematical framework of magnetic inversion, beginning with the forward problem, which establishes the relationship between magnetic susceptibility and observed data. We then discuss the inverse problem, focusing on the strategies used to recover the subsurface model. Finally, we explore the variational and scale-space approaches, highlighting their strengths and limitations in addressing the challenges of magnetic inversion. In particular, we focus on dictionary-based inversion techniques, which we subsequently reformulate into a trainable dictionary framework.
@@ 2.2 Inverse problem
The inverse problem seeks to recover the true subsurface model ${\bf m}$ from observed magnetic data ${\bf d}$, given the forward operator ${\bf A}$, defined in Eq. 2. Due to the ill-posed nature of magnetic inversion, infinitely many solutions satisfy this relationship. Therefore, additional constraints and regularization techniques are necessary to obtain physically plausible and stable solutions.
To address this challenge, theoretical frameworks are employed to incorporate prior knowledge and impose regularization on the solution. Our work focuses on leveraging scale-space methods to tackle the ambiguities inherent in this inverse problem. We demonstrate the effectiveness of these methods by comparing them with established variational approaches, highlighting their potential to improve magnetic data inversion accuracy.
@@
    Analysis-based regularization: To encode prior expectations about the model ${\bf m}$, we introduce a regularization function $R({\bf m})$. Such a prior is also obtained when considering the Maximum A Posteriori (MAP) estimate. It seeks the model ${\bf m}$ that maximizes the posterior, given the observations ${\bf d}$ and the prior, leading to the following optimization problem or loss function [31, 22, Section 3.5.6]:
@@
${\bf z}^{(j+1)}=\sigma_{\tau_{j}}\left({\bf z}^{(j)}-\mu_{j}{\bm{\mathrm{\Psi}}}^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}{\bf z}^{(j)}-\mathbf{d}\right)\right),$ (6)
where $\mu_{j}$ is the step-size, $\tau_{j}$ is the thresholding parameter and ${\bf z}^{(j)}$ represents the value of ${\bf z}$ at $j^{\text{th}}$ iteration. This approach is one of the most popular for solving the optimization problems in Eq. (5) and is thoroughly discussed in [10].
@@
where, $z_{i}$ represents the $i^{\text{th}}$ component of ${\bf z}$. The soft-thresholding operator reduces the magnitude of coefficients by a threshold $\tau$, setting small values to zero, thereby enforcing sparsity and enhancing the stability and robustness of the solution [8, 9]. Here, unlike traditional approaches that iterate to full convergence, we use scale-space methods that terminate early, capturing dominant structures while avoiding overfitting [6]. In our context, we can change $\tau$ as the iteration progresses. Indeed, it is straightforward to see that if we let $\tau_{j}\rightarrow 0$ as $j\rightarrow\infty$, then the only term that is important is the data fitting term, which implies convergence to a solution that fits the data as $j$ grows. Next, we show how to use the scale-space formulation of sparse recovery in the context of dictionary learning for magnetic inversion.
@@ 3 Learning scale space
${\ell}({\bm{\mathrm{\Psi}}})=\frac{1}{2}{\mathbb{E}}_{{\bf m}}\Big{[}\|\widehat{\bf m}-{\bf m}\|^{2}\Big{]}+\frac{1}{2}{\mathbb{E}}_{{\bf d}}\Big{[}\|\widehat{\bf d}-{\bf d}\|^{2}\Big{]}.$ (10)
Hence, the optimal ${\bm{\mathrm{\Psi}}}$ is the one that minimizes the mean loss over all models in our data set as well as the one that fits the data by minimizing over the mean loss over all data.
@@
One standard approach for solving such a problem relies on stochastic gradient descent [5]. In this approach, we sample the model space ${\cal M}$, use the forward modeling to simulate the data and then use the network (6) to recover a solution and evaluate the loss. We then use automatic differentiation to compute the gradient of the loss with respect to ${\bm{\mathrm{\Psi}}}$ and take a step in the direction of negative gradient. The method is summarized in Algorithm 1. While the algorithm presented here is rather generic, we have experimented with two types of dictionary for the scale-space methods, a single shared dictionary and a dynamic iteration-dependent dictionary that is sometimes also referred to as the unrolled dictionary.
Algorithm 1 Dictionary learning
Data =${\cal M}=\{{\bf m}_{1},\ldots,{\bf m}_{N}\}$
while not converge do
     Randomly sample ${\bf m}$ from ${\cal M}$
@@
     Use Eq. (6) to compute $\widehat{\bf m}$ given ${\bf d}$ and use Eq. (9) to compute $\widehat{\bf d}$ from $\widehat{\bf m}$
     Estimate the loss $\ell=\frac{1}{2}\|\widehat{\bf m}-{\bf m}\|^{2}+\frac{1}{2}\|\widehat{\bf d}-{\bf d}\|^{2}$
     Compute $\delta{\bm{\mathrm{\Psi}}}=\nabla_{{\bm{\mathrm{\Psi}}}}\ell$
@@
3.1 Shared ${\bm{\mathrm{\Psi}}}$ dictionary
The optimization problem in Eq. (5) and the corresponding iterative method in Eq. (6) to solve it, uses a single dictionary ${\bm{\mathrm{\Psi}}}$. In the context of neural networks, the dictionary is a convolution that is shared between the different layers. The advantages of such a convolution is simplicity and robustness. Indeed, the method can be easily interpreted as dictionary learning for the $\ell_{1}$ recovery problem previously discussed in [13]. However, while this approach simplifies computations, it limits the model’s ability to adapt to complex and dynamic patterns in the data.
@@ 3.2 Unrolled ${\bm{\mathrm{\Psi}}}$ dictionary
A less restrictive approach uses an iteration-dependent dictionary. This enables it to adapt dynamically throughout the training process, yielding further degrees of freedom. In this case, each layer is assigned a unique dictionary ${\bm{\mathrm{\Psi}}}^{(i)}$, allowing the model to better capture intricate and evolving patterns in the data. This is achieved by using a separate set of weights for each iteration, unlike the fixed-dictionary approach where weights are shared. The disadvantage of this approach is the complete loss of some global objective function which is minimized. Nonetheless, this is exactly the power of scale-space methods which introduce a dynamical process for the regularization without it having to be rooted in optimization. The iterative method for the case of unrolled dictionary can be written as,
${\bf z}^{(j+1)}=\sigma_{\tau}\left({\bf z}^{(j)}-\mu({\bm{\mathrm{\Psi}}}^{(j)})^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}^{(j)}{\bf z}^{(j)}-\mathbf{d}\right)\right).$ (11)
@@ 4 Numerical experiments
In this section, we evaluate experiments in 3D that demonstrate the advantage of learning regularization over pre-determined models. We used synthetic data and experimented with four different methods, (i) a variational approach that is based on the gradients of the model, (ii) a sparse recovery with pre-determined fixed dictionary that is based on the cosine transform, (iii) a shared learned dictionary that approximately minimizes Eq. (5), and (iv) a learned unrolled dictionary that is iteration-dependent (Eq. (11)). The first two approaches were used as baselines to the learned approaches proposed in this work.
@@ 4.1 Dataset
We performed experiments using a $64\times 64\times 32$ uniform grid of $100\text{m}^{3}$ cells with a baseline magnetic susceptibility of $\chi_{v}=0$. A randomly selected number of features were added to the model to form ellipsoid shaped areas of increased magnetic susceptibility with exponential radial decay (see Table 1). Such models simulate the existence of mineral deposits that are commonly present in geological structures such as porphyries [21]. Given these models, we used Algorithm 1 to train the dictionary(ies) ${\bm{\mathrm{\Psi}}}$ with a training set of 2000 samples, a validation set of 100 samples. The model with the lowest loss on the validation set was used for final testing on the unseen test set consisting of 500 samples. We added $1\%$ relative noise to the forward data to simulate measurement errors before performing the inversion.

### v2

@@ Abstract
Magnetic data inversion is an important tool in geophysics, used to infer subsurface magnetic susceptibility distributions from surface magnetic field measurements. This inverse problem is inherently ill-posed, characterized by non-unique solutions, depth ambiguity, and sensitivity to noise. Traditional inversion approaches rely on predefined regularization techniques to stabilize solutions, limiting their adaptability to complex or diverse geological scenarios. In this study, we propose an approach that integrates variable dictionary learning and scale-space methods to address these challenges. Our method employs learned dictionaries, allowing for adaptive representation of complex subsurface features that are difficult to capture with predefined bases. Additionally, we extend classical variational inversion by incorporating multi-scale representations through a scale-space framework, enabling the progressive introduction of structural detail while mitigating overfitting. We implement both fixed and dynamic dictionary learning techniques, with the latter introducing iteration-dependent dictionaries for enhanced flexibility. Using a synthetic dataset to simulate geological scenarios, we demonstrate significant improvements in reconstruction accuracy and robustness compared to conventional variational and dictionary-based methods. Our results highlight the potential of learned dictionaries, especially when coupled with scale-space dynamics, to improve model recovery and noise handling. These findings underscore the promise of our data-driven approach for advance magnetic data inversion and its applications in geophysical exploration, environmental assessment, and mineral prospecting. The code is publicly available at: https://github.com/ahxmeds/magnetic-inversion-dictionary.git.
@@ 1 Introduction
However, magnetic inversion is a classic example of a highly ill-posed inverse problem, where multiple subsurface models can equally explain the same observed magnetic anomalies. Magnetic signals attenuate with increasing depth, thereby limiting the resolution of deeper structures and compounding the ambiguity [28]. Noise in the data and variability in magnetization properties, such as induced and remanent components, can significantly distort the inversion outcomes [11]. To address these challenges, classical inversion methods incorporate regularization, such as $\ell_{1}$-$\ell_{2}$ norms, to balance sparsity and smoothness [15, 28]. Probabilistic frameworks, including linear stochastic inversion introduced by Franklin [12] and refined by Tarantola and Valette [24, 20], initially provided a foundation for addressing inversion ambiguities. Building on this probabilistic approach, Bosch et al. [3, 4] used Monte Carlo methods for gravity inversion to derive posterior probability densities, capturing the range of plausible models consistent with the data. Chasseriau and Chouteau [7] emphasized integrating prior geological information through three-dimensional gravity data inversion constrained by an a priori covariance model. Incorporating spatial correlations of subsurface properties enhances stability and geological realism, overcoming the limitations of traditional deterministic methods. Nonetheless, these conventional approaches rely on pre-determined regularization schemes, which may limit their adaptability to complex geological scenarios. More recently, work in the related field of imaging inverse problems has demonstrated the effectiveness of multiscale convolutional dictionary learning as a method for solving challenging problems such as medical image reconstruction [16].
In this work, we propose a similar approach for magnetic inversion by integrating dictionary learning [13] with a scale-space framework. Our framework aims to learn a dynamic dictionary that evolves over iterations, creating a versatile inversion process that can approximate complex models which fixed dictionaries often fail to represent. Numerical experiments demonstrate this method not only outperforms classical methods in terms of accuracy and robustness but also offers greater flexibility in handling diverse geological structures. This analysis highlights the promise of a scale-space approach as an alternative for standard dictionary learning in geophysical inversion.
@@ 2 Overview of the Problem
Magnetic inversion aims to reconstruct a three-dimensional magnetic susceptibility model from two-dimensional surface magnetic data. According to Gauss’s theorem, infinitely many non-unique source distributions can produce the same observed magnetic field. To address these challenges, inversion methods incorporate prior information and regularization strategies to ensure physically plausible and geologically meaningful solutions.
This section introduces the mathematical framework of magnetic inversion, beginning with the forward problem, which establishes the relationship between magnetic susceptibility and observed data. We then discuss the inverse problem and the strategies used to recover a subsurface model. Finally, we explore the variational and scale-space approaches, highlighting their strengths and limitations in addressing the challenges of magnetic inversion. In particular, we focus on dictionary-based inversion techniques, which we subsequently reformulate into a trainable dictionary framework.
@@ 2.2 Inverse Problem
The inverse problem seeks to recover the true subsurface model ${\bf m}$ from observed magnetic data ${\bf d}$, given the forward operator ${\bf A}$, defined in Eq. 2. Due to the ill-posed nature of magnetic inversion, additional constraints and regularization techniques incorporating prior knowledge are necessary to obtain physically plausible and stable solutions. Our work focuses on leveraging scale-space methods to tackle the ambiguities inherent in this inverse problem. We demonstrate the effectiveness of these methods by comparing them with established variational approaches, highlighting their potential to improve accuracy.
@@
    Analysis-based regularization: To encode prior expectations about the model ${\bf m}$, we introduce a regularization function $R({\bf m})$. Such a prior is also obtained when considering the Maximum A Posteriori (MAP) estimate. It seeks the model ${\bf m}$ that maximizes the posterior, given the observations ${\bf d}$ and the prior, leading to the following optimization problem or loss function [32, 23, Section 3.5.6]:
@@
${\bf z}^{(j+1)}=\sigma_{\tau}\left({\bf z}^{(j)}-\mu_{j}{\bm{\mathrm{\Psi}}}^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}{\bf z}^{(j)}-\mathbf{d}\right)\right),$ (6)
where $\mu_{j}$ is the step-size, ${\bf z}^{(j)}$ represents the value of ${\bf z}$ at $j^{\text{th}}$ iteration and $\tau$ is the thresholding parameter. This approach is one of the most popular for solving the optimization problems in Eq. (5) and is thoroughly discussed in [10].
@@
where, $z_{i}$ represents the $i^{\text{th}}$ component of ${\bf z}$. The soft-thresholding operator reduces the magnitude of coefficients by a threshold $\tau$, setting small values to zero, thereby enforcing sparsity and enhancing the stability and robustness of the solution [8, 9]. Here, unlike traditional approaches that iterate to full convergence, we use scale-space methods that terminate early, capturing dominant structures while avoiding overfitting [6]. Next, we show how to use the scale-space formulation of sparse recovery in the context of dictionary learning for magnetic inversion.
@@ 3 Learning Scale Space
${\ell}({\bm{\mathrm{\Psi}}})=\frac{1}{2}{\mathbb{E}}_{{\bf m}}\Big{[}\|\widehat{\bf m}-{\bf m}\|^{2}\Big{]}+\alpha\frac{1}{2}{\mathbb{E}}_{{\bf d}}\Big{[}\|\widehat{\bf d}-{\bf d}\|^{2}\Big{]},$ (10)
where $\alpha$ is a hyper-parameter that weighs the contribution of the forward-data misfit term relative to the dictionary-based model misfit. The additional loss term on the forward data promotes solutions that match the observed data from the subset of possible $\widehat{\bf m}$ which may have the same $\ell_{2}$-norm distance from the true model. In our study, we used $\alpha=1$ for convenience. Hence, the optimal ${\bm{\mathrm{\Psi}}}$ is the one that minimizes the mean loss over all models in our data set as well as the one that fits to the observed data by minimizing over the mean loss over all data.
@@
One standard approach for solving such a problem relies on stochastic gradient descent [5]. In this approach, we sample the model space ${\cal M}$, use the forward modeling to simulate the data and then use the network (6) to recover a solution and evaluate the loss. We then use automatic differentiation to compute the gradient of the loss with respect to ${\bm{\mathrm{\Psi}}}$ and take a step in the direction of negative gradient. The method and algorithm is summarized in subsections 3.1 and 3.2 below where we have experimented with two types of dictionary for the scale-space methods: a single shared dictionary and a dynamic iteration-dependent dictionary that is sometimes also referred to as the unrolled dictionary.
3.1 Shared ${\bm{\mathrm{\Psi}}}$ Dictionary
The optimization problem in Eq. (5) and the corresponding iterative method in Eq. (6) to solve it, uses a single dictionary ${\bm{\mathrm{\Psi}}}$, shown in Algorithm 1. In the context of neural networks, the dictionary is a convolution that is shared between the different layers. The advantages of such a convolution is simplicity and robustness. Indeed, the method can be easily interpreted as dictionary learning for the $\ell_{1}$ recovery problem previously discussed in [13]. However, while this approach simplifies computations, it limits the model’s ability to adapt to complex and dynamic patterns in the data.
Algorithm 1 Single Shared Dictionary Learning
Data =$\mathcal{M}=\{{\bf m}_{1},\ldots,{\bf m}_{N}\}$, operator ${\bf A}$, learning rate $\alpha$, iterations $N$
Initialize a single dictionary ${\bm{\mathrm{\Psi}}}$ of size $1\times C\times H\times W\times D$
while not converged do
     Randomly sample ${\bf m}$ from $\mathcal{M}$
@@ 3.1 Shared ${\bm{\mathrm{\Psi}}}$ Dictionary
     ${\bf z}^{(0)}\leftarrow\mathbf{0}$triangleright Initialize latent representation
     for $j=0$ to $N-1$ do
         ${\bf z}^{(j+1)}=\sigma_{\tau}\left({\bf z}^{(j)}-\mu_{j}{\bm{\mathrm{\Psi}}}^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}{\bf z}^{(j)}-{\bf d}\right)\right)$triangleright Soft-thresholding Eq. (6)
     end for
     Compute $\widehat{{\bf m}}={\bm{\mathrm{\Psi}}}\,{\bf z}^{(N)}$ and $\widehat{{\bf d}}={\bf A}\,\widehat{{\bf m}}$triangleright Reconstruction step Eq. (9)
     Estimate the loss $\ell=\frac{1}{2}\|\widehat{{\bf m}}-{\bf m}\|^{2}+\frac{1}{2}\|\widehat{{\bf d}}-{\bf d}\|^{2}$triangleright Loss function Eq. (10)
     Compute $\delta{\bm{\mathrm{\Psi}}}=\nabla_{{\bm{\mathrm{\Psi}}}}\,\ell$triangleright Automatic differentiation
@@ 3.2 Unrolled ${\bm{\mathrm{\Psi}}}$ Dictionary
A less restrictive approach uses an iteration-dependent dictionary, shown in Algorithm 2. This allows the dictionary to adapt dynamically throughout the training process, yielding further degrees of freedom. In this case, each layer is assigned its own dictionary ${\bm{\mathrm{\Psi}}}_{(j)}$, allowing the model to better capture intricate and evolving patterns in the data. This is achieved by using a separate set of weights for each iteration, unlike the fixed-dictionary approach where weights are shared. However, because each iteration has its own dictionary, we lose a single global objective in the strict variational sense - there is no single functional whose minimizer the algorithm is guaranteed to converge to. Nonetheless, this is exactly the power of scale-space methods which introduce a dynamical process for the regularization without it having to be rooted in optimization. The iterative method for the case of unrolled dictionary can be written as,
${\bf z}^{(j+1)}=\sigma_{\tau}\left({\bf z}^{(j)}-\mu_{j}{\bm{\mathrm{\Psi}}}_{(j)}^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}_{(j)}{\bf z}^{(j)}-\mathbf{d}\right)\right).$ (11)
@@
Algorithm 2 Unrolled Dictionary Learning
Data =$\mathcal{M}=\{{\bf m}_{1},\ldots,{\bf m}_{N}\}$, operator ${\bf A}$, learning rate $\alpha$, iterations $N$
Initialize a set of dictionaries ${\bm{\mathrm{\Psi}}}_{(1)},...,{\bm{\mathrm{\Psi}}}_{(N)}$ of size $N\times C\times H\times W\times D$
while not converged do
     Randomly sample ${\bf m}$ from $\mathcal{M}$
     Compute ${\bf d}={\bf A}{\bf m}+\epsilon$
     ${\bf z}^{(0)}\leftarrow\mathbf{0}$triangleright Initialize latent representation
     for $j=0$ to $N-1$ do
         ${\bf z}^{(j+1)}=\sigma_{\tau}\left({\bf z}^{(j)}-\mu_{j}{\bm{\mathrm{\Psi}}}_{(j)}^{\top}{\bf A}^{\top}\left({\bf A}{\bm{\mathrm{\Psi}}}_{(j)}{\bf z}^{(j)}-\mathbf{d}\right)\right)$triangleright Soft-thresholding Eq. (11)
     end for
     Compute $\widehat{{\bf m}}={\bm{\mathrm{\Psi}}}\,{\bf z}^{(N)}$ and $\widehat{{\bf d}}={\bf A}\,\widehat{{\bf m}}$triangleright Reconstruction step Eq. (9)
     Estimate the loss $\ell=\frac{1}{2}\|\widehat{{\bf m}}-{\bf m}\|^{2}+\frac{1}{2}\|\widehat{{\bf d}}-{\bf d}\|^{2}$triangleright Loss function Eq. (10)
     Compute $\delta{\bm{\mathrm{\Psi}}}_{(j)}=\nabla_{{\bm{\mathrm{\Psi}}}_{(j)}}\,\ell$ for all $j$
     Update ${\bm{\mathrm{\Psi}}}_{(j)}\leftarrow{\bm{\mathrm{\Psi}}}_{(j)}-\alpha\delta{\bm{\mathrm{\Psi}}}_{(j)}$ for all $j$
end while
@@ 4 Numerical Experiments
In this section, we evaluate experiments in 3D that demonstrate the advantage of learning regularization over pre-determined models. We used synthetic data and experimented with four different methods, (i) a variational approach that is based on the gradients of the model, (ii) a sparse recovery with pre-determined fixed dictionary that is based on the cosine transform, (iii) a shared learned dictionary that approximately minimizes Eq. (5), and (iv) a learned unrolled dictionary that is iteration-dependent (Eq. (11)). The first two approaches were used as baselines to the learned approaches proposed in this work. The code to reproduce our results is publicly available at: https://github.com/ahxmeds/magnetic-inversion-dictionary.git.
@@ 4.1 Dataset
We performed experiments using a $64\times 64\times 32$ uniform grid of $100\text{m}^{3}$ cells with a baseline magnetic susceptibility of $\chi_{v}=0$. A randomly selected number of features were added to the model to form ellipsoid shaped areas of increased magnetic susceptibility with exponential radial decay (see Table 1). Such models simulate the existence of mineral deposits that are commonly present in geological structures such as porphyries [22]. Given these models, we used Algorithm LABEL:alg1 to train the dictionary(ies) ${\bm{\mathrm{\Psi}}}$ with a training set of 2000 samples, a validation set of 100 samples. The model with the lowest loss on the validation set was used for final testing on the unseen test set consisting of 500 samples. We added $1\%$ relative noise to the forward data to simulate measurement errors before performing the inversion.